find the roots using simple fixed point. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

2x^5+x^4-2x-1=0
1
Expert's answer
2020-10-26T18:55:50-0400
"2x^5+x^4-2x-1=x^4(2x+1)-(2x+1)="

"=(2x+1)(x^4-1)=(2x+1)(x^2-1)(x^2+1)="

"=(2x+1)(x+1)(x-1)(x^2+1)"

"x1=-1, x2=-\dfrac{1}{2}, x3=1"

We know that there is a solution for the equation "2x^5+x^4-2x-1=0" in "[0.5,1.5]."


"x=\big(\dfrac{-x^4+2x+1}{2}\big)^{1/5}"

"x_{i+1}=\big(\dfrac{-x_i^4+2x_i+1}{2}\big)^{1/5}"

"\begin{matrix}
 i & x_i \\
 0 & 0.5\\
 1 & 0.993670 \\
 2 & 1.001239\\
 3 & 0.999751\\
 4 & 1.000050\\
 5 & 0.999990\\
 6 & 1.000002\\
 7 &1.000000 \\
 8 & 1.000000\\
 
\end{matrix}"


"\varepsilon =\big|\dfrac{n_{i+1}-n_i}{n_i}\big|\cdot 100\%"

"\varepsilon =\big|\dfrac{0.993670-0.5}{0.5}\big|\cdot 100\%=98.734\%"

"\varepsilon =\big|\dfrac{1.001239-0.993670}{0.993670}\big|\cdot 100\%=0.7617\%"

"\varepsilon =\big|\dfrac{0.999751-1.001239}{1.001239}\big|\cdot 100\%=0.1486\%"

"\varepsilon =\big|\dfrac{1.000050-0.999751}{0.999751}\big|\cdot 100\%=0.0299\%"

"\varepsilon =\big|\dfrac{0.999990-1.000050}{1.000050}\big|\cdot 100\%=0.0060\%"

"\varepsilon =\big|\dfrac{0.999998-0.999990}{0.999990}\big|\cdot 100\%=0.0016\%"

"\varepsilon =\big|\dfrac{1.000002-0.999998}{0.999998}\big|\cdot 100\%=0.0012\%"

"\varepsilon =\big|\dfrac{1.000000-1.000002}{1.000002}\big|\cdot 100\%=0.0002\%"

If "\varepsilon\leq0.05\%," then "x=0.999998"


The root is "1.00000"